The sum of the digits of a two digit number is 16. If the digits were reversed, and the new number were added to the original number the sum would be 176. Find the original number.
Thanxs sooo much

Answer by Nate(3500)

Digit: |x|v|
x + v = 16
10v + x + 10x + v = 176
Combine:
11x + 11v = 176
+
-11x - 11v = -176
=
0 = 0
The digit is: 88